OpenStudy (anonymous):

One of the steps in the extraction of iron from its ore is the reduction of ironII oxode by CO at 900 degrees celsius FEO(s)+CO(g)--><-- Fe(s)+CO2(g)

OpenStudy (anonymous):

IF CO is allowed to react with an excess of FEO, calculate the mole fractions of CO and CO2 at equilibrium . State any assumptions

OpenStudy (anonymous):

i'll be giving a medal :)))

OpenStudy (aaronq):

do you have any other info? like the eq. constant?

OpenStudy (anonymous):

so at the equilibrium we establish that dG(T)=0... so the chemical equilibrium would be \[DeltaG ^{o}=-RTln K\] that in this case K=[CO2]/[CO] (as the activities of solids are equal to )

OpenStudy (anonymous):

also the amount of moles that will react in CO will be equal to the moles of CO2... we will establish:\[DeltaG ^{o}=-RTln \left(\begin{matrix}\frac{ X }{ 1-X} \end{matrix}\right)\] 1 because we're gonna assume an initial mole fraction of 1 for CO

OpenStudy (anonymous):

so, you need to calculate dG by hess' law... substitute and thats it

OpenStudy (anonymous):

No, you need to get dG by Hess' law, and then get it at the given temperature the formula for getting the dG at (1173 K) would be:\[\frac{ dG ^{o}_{T} }{ T }=\frac{ dG ^{o}_{298} }{ 298 }-\int\limits_{298}^{T}\frac{ dH ^{o}_{T} }{ T ^{2} }\]

OpenStudy (anonymous):

mmm you would also need\[dH ^{o}_{T}=dH ^{o}_{298}+\int\limits_{298}^{T}\Delta(nCp)dT\]

OpenStudy (anonymous):

in case your Cp is variable (you'd also use Hess' law here), integrate and dont substitute the T value... just use it in the dG equation and the integrate again

OpenStudy (anonymous):

@martaamador62 all clear?

OpenStudy (anonymous):

@emilykrucker yes thanks for helping me :)